An electrician needs 4 3/4 rolls of electrical wire to wire each room in a house. How many rooms can he wire with 9 1/2 rolls of wire?
Answer by addingup(3677)

9 1/2 divided by 4 3/4, no?
(9 1/2)/(4 3/4)
19/2)/19/4) Divide:
19*4/2*19= 76/38= 2 rooms